OpenStudy (rjtrain1):

The solution to the equation x + 11 = 18 is x = 7. True False

OpenStudy (awesomedude2078):

C. hope this helps :)

OpenStudy (awesomedude2078):

because 8 (m+11) = 8*m + 8 * 11 = 8m + 88

OpenStudy (rjtrain1):

thank you all who answers

OpenStudy (awesomedude2078):

you're welcome ;)

OpenStudy (awesomedude2078):

i'd say A

OpenStudy (rjtrain1):

and stick here please I will keep changing the question

OpenStudy (rjtrain1):

when I close the question I will be done with this 6th grade quiz

OpenStudy (rjtrain1):

btw

OpenStudy (rjtrain1):

and can you keep replying to the answers